OpenStudy (anonymous):

help me solve for x; 5/8x+1/16x=11/16+x

OpenStudy (mertsj):

\[\frac{5}{8}x+\frac{1}{16}x-x=\frac{11}{16}\]

OpenStudy (theviper):

\[5/8x+1/16x=11/16x\] =>\[5/8x+1\times1/16x=11/16+x\]

OpenStudy (anonymous):

so what is x

OpenStudy (mertsj):

\[(\frac{5}{8}+\frac{1}{16}-1)x=\frac{11}{16}\]

OpenStudy (mertsj):

\[\frac{-5}{16}x=\frac{11}{16}\]

OpenStudy (mertsj):

\[\frac{-16}{5}\times\frac{-5}{16}x=\frac{11}{16}\times\frac{-16}{5}\]

OpenStudy (mertsj):

\[x=\frac{-11}{5}\]